OpenStudy (anonymous):

Which of the following are solutions to the equation below? Check all that apply. 12x2 - 5x - 1 = 1 A. -2 B. -1 C. -2/3 D. -1/4 E. 1/4 F. 2/3

OpenStudy (anonymous):

@nathaly95 this same like last one:)

OpenStudy (anonymous):

D,F :)

OpenStudy (anonymous):

\[\Delta=b^2-4ac\] \[\Delta=(-5)^2-4*12*(-2)\]=121 \[\sqrt{\Delta}=11\]

OpenStudy (anonymous):

D. -1/4 is correct

OpenStudy (anonymous):

X1=(5-11)/24 = -1/4 X2=(5+11)/24=16/24=2/3

OpenStudy (anonymous):

@nathaly95 you will have to solve it now

OpenStudy (anonymous):

ok